OpenStudy (anonymous):

solve the given differential equation by undetermined coefficients y'''+y"-4y'+4y=5-e^x+e^2x

OpenStudy (anonymous):

im sorry im not sure how to do this but couldnt you combine like terms?

myininaya (myininaya):

You first to find the homogeneous solution.

myininaya (myininaya):

You can do that pretty easy if you know how to solve: \[r^3+r^2-4r+4=0\]

myininaya (myininaya):

That can be done by factoring by grouping.

myininaya (myininaya):

Let me know if and when you have the homogeneous solution, then you can proceed to find the particular solution by the method in your post.

OpenStudy (anonymous):

Ok for my yc I got yc =c1e^-2x+c2e^2x+c3e^x

OpenStudy (anonymous):

Is that what you got

myininaya (myininaya):

looks awesome! :)

myininaya (myininaya):

now notice on the left hand side of your differential equation you have a constant+constant*e^x+constant*e^(2x)

OpenStudy (anonymous):

Ok

myininaya (myininaya):

So we are going to take your homogeneous solution and add it to something that looks like that so what i'm trying to say we are going to try the following as a solution: \[y=c_1e^{-2x}+c_2e^{2x}+c_3e^{x}+c_4 \]

OpenStudy (anonymous):

So would my yp be yp=a+bxe^x+cxe^2x

OpenStudy (anonymous):

Ok

myininaya (myininaya):

So we have to find y' y'' y''' and plug all of those in to find our constants

OpenStudy (anonymous):

Ok

myininaya (myininaya):

http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx I really believe this site will help you to get the feel for what solutions to try. But sometimes it is trial and error.

OpenStudy (anonymous):

Ok but are you still going to help me

myininaya (myininaya):

Yes. I was just giving the site to you because I thought it would be helpful to you to determine what things to try in the future.

myininaya (myininaya):

Just let me know what else you need help on.

OpenStudy (anonymous):

Ok so the what did you got for the final answer

myininaya (myininaya):

have you plug in y^(3), y^(2),y^(1), and y yet and then grouped your liked terms together and compared to the other sides to find out what the constant coefficients should be

OpenStudy (anonymous):

Yes and my remainder after was -4be^x-4e^2x+4a

OpenStudy (anonymous):

But idk what to do from there

myininaya (myininaya):

\[y=c_1e^{-2x}+c_2e^{2x}+c_3e^{x}+c_4 \\ y'=-2c_1e^{-2x}+2c_2e^{2x}+c_3e^{x}+0\] then so on with the other derivatives c_4 should be easy to find

myininaya (myininaya):

like you don't have any other constant terms in any of the other derivatives so we know 4c_4=5

myininaya (myininaya):

show me the second and third derivative and I will show you all you need to do is plug in and collect likes and compare sides to get the constants

OpenStudy (anonymous):

It's going to take a whil to typ can you just show me step by step of what you did

myininaya (myininaya):

2 and 3rd derivative only have three terms

myininaya (myininaya):

I will find the 2nd one for you but i want you to show me you have put some work into problem

myininaya (myininaya):

\[y^{(2)}=4c_1e^{-2x}+4c_2e^{2x}+c_3e^{x}\]

OpenStudy (anonymous):

I have I'm just really lost bc of the remainder I got after lugging them all in

myininaya (myininaya):

So what did you get the third derivative?

OpenStudy (anonymous):

y"'=-8c1e^-2x+8c2e^2x+c3e^x

myininaya (myininaya):

Great. I basically just wanted to make sure you weren't having problems differentiating and clearly that isn't your trouble here. So next part lets plug in.

OpenStudy (anonymous):

Ok

myininaya (myininaya):

\[y^{(3)}+y^{(2)}-4y^{(1)}+4y \\ =(-8c_1e^{-2x}+8c_2e^{2x}+c_3e^{x}) \\ +(4c_1e^{-2x}+4c_2e^{2x}+c_3e^{x}) \\ -4(-2c_1e^{-2x}+2c_2e^{2x}+c_3e^{x}) \\ +4(c_1e^{-2x}+c_2e^{2x}+c_3e^{x}+c_4) \] And this side is suppose to equal the other side \[1e^{2x}-1e^{x}+5\]

myininaya (myininaya):

we only have one constant term on both sides right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So what is next

myininaya (myininaya):

so we already know that 4c_4=5 which I already said now let's look at the e^x terms

myininaya (myininaya):

group all the e^x terms together on left hand side of the equation

OpenStudy (anonymous):

So c = 1/5

myininaya (myininaya):

like this: \[c_3e^{x}+c_3e^{x}-4c_3e^{x}+4c_3e^{x}=-1e^{x} \\ \text{ \implies } c_3+c_3-4c_3+4c_3=-1 \]

myininaya (myininaya):

well c_4=5/4

myininaya (myininaya):

you divide both sides by 4 to solve for c_4

OpenStudy (anonymous):

Ok my by hahahand what is the a and b

myininaya (myininaya):

what is a and b?

myininaya (myininaya):

have you found c_3 yet?

myininaya (myininaya):

Do you see how I got the equation involving the constant c_3?

OpenStudy (anonymous):

Yes but I'm to nurse about it I got would it be. 1

myininaya (myininaya):

Ok we have \[c_3e^{x}+c_3e^{x}-4c_3e^{x}+4c_3e^{x}=-1e^{x} \\ \text{ \implies } c_3+c_3-4c_3+4c_3=-1 \] I know 1+1-4+4=2+0=2 so I know we have \[2c_3=-1 \]

myininaya (myininaya):

I know to solve for c_3 I must divide both sides by 2

OpenStudy (anonymous):

So it's 1/2

myininaya (myininaya):

yes all we are doing is just solving a bunch on linear equations (like the ones from algebra)

myininaya (myininaya):

You try to find the constant we named c_2

OpenStudy (anonymous):

Ok so c2 would be 1/4

myininaya (myininaya):

\[y^{(3)}+y^{(2)}-4y^{(1)}+4y \\ =(-8c_1e^{-2x}+8c_2e^{2x}+c_3e^{x}) \\ +(4c_1e^{-2x}+4c_2e^{2x}+c_3e^{x}) \\ -4(-2c_1e^{-2x}+2c_2e^{2x}+c_3e^{x}) \\ +4(c_1e^{-2x}+c_2e^{2x}+c_3e^{x}+c_4) \\ =1e^{2x}-1e^{x}+5 \] rememeber we need to compare both sides of this above: gather all the e^{2x}'s on the left hand side just as I did before. so first line we have 8c_2e^(2x) second line we have +4c_2e^{2x) third line we have -4(2c_2e^{2x)) fourth line we have +4(c_2e^(2x)) and on the other side we have 1e^(2x) so we have know we must have the equation \[8c_2+4c_2-8c_2+4c_2=1 \]

myininaya (myininaya):

Is this what you tried to solve?

OpenStudy (anonymous):

So. 1/8

myininaya (myininaya):

yes :)

myininaya (myininaya):

now last constant c_1

OpenStudy (anonymous):

Ci is 1/8 as wel

myininaya (myininaya):

be careful with that

myininaya (myininaya):

do we have any e^(-2x) 's on the other side?

myininaya (myininaya):

\[(-8c_1+4c_1-8c_1+4c_1)e^{-2x}=0e^{-2x} \\ 8c_1=0 \] We did not did we ?

myininaya (myininaya):

remember we have 5-e^x+e^(2x)

myininaya (myininaya):

that didn't consist of any e^{-2x) terms so here we find out we didn't even need to include the c_1e^(-2x) in our solution

myininaya (myininaya):

we chose this as our solution and we found the constants\[y=c_1e^{-2x}+c_2e^{2x}+c_3e^{x}+c_4 \] so last step is to replace those constants with what we found

OpenStudy (anonymous):

Son 0

myininaya (myininaya):

yes c1 is 0

myininaya (myininaya):

\[y=c_2e^{2x}+c_3e^{x}+c_4 \] but you did find constant values for these other c's that weren't 0

OpenStudy (anonymous):

moos the final set Iwould e wh.

OpenStudy (anonymous):

So what wod the yin Laing look like and please Jung tell me c Inge do work on home therhtuff